Given that 0.101mol of an acid H2X weighs 12.9g .calculate (a) molar mass of the acid (b) :relative atomic mass of X
1
Expert's answer
2021-10-27T06:45:10-0400

a) n = m / M

M = m / n = 12.9 / 0.101 = 127.72 g/mol


b) M (H2X) = 127.72 g/mol

M (H) = 1 g/mol

M (X) = 127.72 - (1 x 2) = 125.72 g/mol